OpenStudy (anonymous):

simplify the monomial: (x6 y3) (5xy2) (-7x5) all of the big numbers at the back are exponents. please help!!!!

OpenStudy (anonymous):

hi goody,, the formula for this is like a^m a^n = a^(m+n) try it

OpenStudy (anonymous):

Simplify. (x4 y3)(4xy2)(−7x3)

OpenStudy (anonymous):

(a^m a^n)b^p b^q= [a^(m+n)][b^()p+q]

OpenStudy (mertsj):

\[4(-7)(x^4)(x)(x^3)(y^3)(y^2)=-28x^8y^5\]

OpenStudy (anonymous):

do you have the idea now how to get it?

OpenStudy (anonymous):

just add the powers of liked terms

OpenStudy (anonymous):

yeah now I sort of get it. Thank you!!!!!

OpenStudy (anonymous):

ok yw